OpenStudy (anonymous):

Find the integral of (1/2)(4/(1+sin(theta)))^2 d*theta from 0 to pi.

myininaya (myininaya):

looks like youmight need tan(theta/2)=t

OpenStudy (anonymous):

Can you please show the work?

myininaya (myininaya):

build a right triangle for the sub I gave you

OpenStudy (anonymous):

So we use u-substitution for u=tan^-1 t?

myininaya (myininaya):

well i was using tan(theta/2)=t

OpenStudy (anonymous):

I don't know how you got that in the first place.

myininaya (myininaya):

also do you know your integral is improper?

OpenStudy (anonymous):

Is it? How so?

myininaya (myininaya):

wait whole it is to pi sorry i thought it said 2pi

myininaya (myininaya):

my substitution is commonly used to write trig functions as rationals http://en.wikipedia.org/wiki/Tangent_half-angle_substitution

myininaya (myininaya):

but the limits are weird tan(0/2)=0=t tan(t/2) goes to inf as t goes to pi so this actually transforms it to an improper integral

myininaya (myininaya):

an improper integral that will converge

myininaya (myininaya):

let me know if you need any further help

myininaya (myininaya):

tan(0/2)=0=t tan(t/2) goes to inf as theta goes to pi *

OpenStudy (anonymous):

\[\frac{1}{2}\int_0^\pi\frac{4}{(1+\sin\theta)^2}~d\theta\] There are at least two approaches to this sort of integral: the Weierstrass substitution that @myininaya provided the link for, and a method involving complex contour integration. Is this calculus or complex analysis?

OpenStudy (anonymous):

This is Calculus 1.

OpenStudy (anonymous):

And the 4 should be squared, it's (4/(1+sin theta))^2

OpenStudy (anonymous):

Oh, in that case it might just be a matter of using some trig identities. It might not be the best way though. \[\frac{1}{2}\int_0^\pi\left(\frac{4}{1+\sin\theta}\right)^2~d\theta\\ \frac{1}{2}\int_0^\pi16\left(\frac{1-\sin\theta}{(1+\sin\theta)(1-\sin\theta)}\right)^2~d\theta~~~~~\text{Note: this introduces a discontinuity}\\ 8\int_0^\pi\left(\frac{1-\sin\theta}{1-\sin^2\theta}\right)^2~d\theta\\ 8\int_0^\pi\left(\frac{1-\sin\theta}{\cos^2\theta}\right)^2~d\theta\\ 8\int_0^\pi\frac{1-2\sin\theta+\sin^2\theta}{\cos^4\theta}~d\theta\\ 8\int_0^\pi\left(\sec^2\theta-\frac{2\sin\theta}{\cos^4\theta}+\frac{\sin^2\theta}{\cos^4\theta}\right)~d\theta\] The antiderivative for the first term is simply \(\tan\theta\), though you can leave it alone for now (you'll see why in a moment). The second requires a substitution, \(u=\cos\theta\). The third involves some intermediate steps before you can use a substitution: \[\frac{\sin^2\theta}{\cos^4\theta}=\frac{1-\cos^2\theta}{\cos^4\theta}=\sec^4\theta-\sec^2\theta\] This gives you \[8\int_0^\pi\left(\sec^2\theta-\frac{2\sin\theta}{\cos^4\theta}+\sec^4\theta-\sec^2\theta\right)~d\theta\\ 8\int_0^\pi\left(\sec^4\theta-\frac{2\sin\theta}{\cos^4\theta}\right)~d\theta\] The first term, which I said you can ignore, disappears. \[8\int_0^\pi\sec^4\theta~d\theta-16\int_0^\pi\frac{\sin\theta}{\cos^4\theta}~d\theta\]

OpenStudy (anonymous):

Thanks a lot.